# SimLTD: Simple Supervised and Semi-Supervised Long-Tailed Object Detection

### v1

@@ Abstract
Recent years have witnessed tremendous advances on modern visual recognition systems. Despite such progress, many vision models still struggle with the open problem of learning from few exemplars. This paper focuses on the task of object detection in the setting where object classes follow a natural long-tailed distribution. Existing approaches to long-tailed detection resort to external ImageNet labels to augment the low-shot training instances. However, such dependency on a large labeled database is impractical and has limited utility in realistic scenarios. We propose a more versatile approach to leverage optional unlabeled images, which are easy to collect without the burden of human annotations. Our SimLTD framework is straightforward and intuitive, and consists of three simple steps: (1) pre-training on abundant head classes; (2) transfer learning on scarce tail classes; and (3) fine-tuning on a sampled set of both head and tail classes. Our approach can be viewed as an improved head-to-tail model transfer paradigm without the added complexities of meta-learning or knowledge distillation, as was required in past research. By harnessing supplementary unlabeled images, without extra image labels, SimLTD establishes new record results on the challenging LVIS v1 benchmark across both supervised and semi-supervised settings.
@@ 1 Introduction
Another interesting direction is to leverage external data to augment the training instances in LVIS. The intuition is that while the rare objects may not appear frequently in natural scenes, they can be found in abundance from object-centric sources such as ImageNet [9]. By training on the union of scene-centric LVIS images and object-centric ImageNet examples, these methods [45, 50, 27] overcome the skewed class distribution by enriching the tail instances with additional whole-image labels. Although appealing, this approach places a hard requirement on having access to a large database of 14M labeled images across 21K object classes, an assumption not necessarily suitable for many practical settings outside of LVIS.
@@
We propose a super simple framework, aptly named SimLTD, to answer this question. We deconstruct the LTD problem into three stages consisting of (i) pre-training on head classes, (ii) transfer learning on tail classes, and (iii) fine-tuning on a small sampled set composed of both head and tail classes. Unlike previous research on multi-stage training [19, 12], our framework allows for the optional use of unlabeled images to further boost LTD. We learn with supplementary unlabeled data in a semi-supervised manner via pseudo-labeling, and therefore do not explicitly rely on any additional instance-level or image-level supervision.
Our work addresses several challenges associated with the head-to-tail model transfer paradigm [18, 39]. For one, we find that the vanilla transfer of data-rich head representations to data-scarce tail classes does not yield sufficient performance improvement because the distribution of head classes is also skewed. Moreover, the naïve application of unlabeled data can aggravate the model’s inductive bias because the unlabeled samples may follow a similar long-tailed distribution. In such case, the trained model will tend to generate more pseudo labels for the head classes, resulting in a larger degree of pseudo-class imbalance. We overcome these obstacles by incorporating data augmentations specifically designed to mitigate class imbalances in both head and tail training stages. Extensive experiments in §4.3 show that stronger pre-trained models on head classes, with and without unlabeled images, transfer well to the tail classes, leading to a desired solution for long-tailed detection.
@@ 2 Related Work
Leveraging Extra Data for LTD There is an emerging trend of using the abundant weak-supervision of ImageNet labels to solve the LVIS problem. Although LVIS is a detection dataset of natural scenes and ImageNet is a classification dataset of object images, they both share 997 overlapping classes between their vocabularies, the intersection of which provides a rich source of 1.5M extra images to sample additional LVIS instances. To effectively learn from a mixed LVIS-ImageNet dataset with a domain gap, MosaicOS [45] leverages mosaic augmentation [1] whereas Detic [50] and RichSem [27] rely on the CLIP [30] classifier to map the semantic concepts between LVIS and ImageNet targets.
@@ 3 The SimLTD Framework
As illustrated in Figure 2, given a long-tailed dataset with categories, we split it into two disjoint subsets: with frequent and common categories appearing in images and with rare classes appearing in images. Furthermore, we have access to an unlabeled dataset of unknown class distribution. Our goal is to use a combination of labeled and unlabeled images to train a unified model optimized for accuracy on a test set comprising both classes in .
Our SimLTD framework consists of three easy steps: (i) representation learning on , (ii) transfer learning on , and (iii) fine-tuning on , a reduced dataset composed of instances per class randomly sampled from . Note that , , and are all still imbalanced, but not as severe as the original long-tailed . We leverage optional unlabeled images in both Steps 1 and 2 but do not explicitly need them for effective LTD. Indeed, experiments in §4.3 show that our fully supervised baselines still exhibit excellent performance and scalability without unlabeled data. A diagram of SimLTD is depicted in Figure 3.
@@ 3.1 The Devil is in the
Discussion We observe a clear trend indicating stronger pre-trained representations, as measured by the AP score on COCO, generally lead to improved rare class detection. This result is both intriguing and encouraging since the models were pre-trained on COCO, a dataset of different scope and size than LVIS. Training from scratch brings out the worst performance with half of the accuracy. The implication of this simple experiment is two-fold: (1) we corroborate prior studies by showing low-shot learning can be improved with transferred representations; and (2) our framework opens up opportunities to self-supervised, semi-supervised, and multi-modal learning, all of which have demonstrated significantly better performance than supervised pre-training. Motivated by these insights, we propose to learn powerful representations on and transfer them to for long-tailed detection. While the past attempts at head-to-tail model transfer [39, 19, 12] could only work by incorporating an extra module for meta-learning or knowledge distillation, we now describe our three-step approach to accomplish this goal without the added complexities.
@@ 3.2 Step 1: Representation Learning on
As summarized in Table 1, we leverage well-known data augmentations to train , which include random image resizing, horizontal flipping, and photometric distortion. We also combine Simple Copy-Paste (SCP) [14] together with Repeat Factor Sampling (RFS) [15] to combat the class imbalance in at both the image and instance levels, analogous to prior research on the importance of image and instance resampling for LTD [37, 42]. Taking a step further to learn even stronger representations on , we propose to train in a semi-supervised manner with unlabeled images by way of pseudo-labeling, which is currently leading the literature on semi-supervised detection. We explore three successful methods of increasing effectiveness to advance semi-supervised representation learning on : SoftER Teacher [35], MixTeacher [26], and MixPL [6].
@@
BAGS [CVPR’20] [22] | N/A | Faster R-CNN | R101-FPN | – | 27.6 | 18.7 | 26.5 | 32.6 |
Seesaw Loss [CVPR’21] [36] | Faster R-CNN | R101-FPN | – | 27.8 | 18.7 | 27.0 | 32.8 | |
@@
ECM [ECCV’22] [7] | Mask R-CNN | R101-FPN | – | 29.4 | 21.9 | 27.9 | 32.3 | |
LogN [IJCV’24] [48] | Mask R-CNN | R101-FPN | – | 29.8 | 22.9 | 28.8 | 31.8 | |
SimLTD Supervised Baseline | Faster R-CNN | R101-FPN | 83 | 31.7 | 24.3 | 32.1 | 34.6 | |
SimLTD Supervised Baseline | Deformable DETR | R101 | 124 | 37.0 | 31.9 | 36.6 | 39.6 | |
@@
SimLTD Supervised Baseline | Faster R-CNN | R50-FPN | 81 | 29.0 | 20.9 | 29.7 | 31.9 | |
SimLTD Supervised Baseline | Deformable DETR | R50 | 115 | 35.0 | 32.0 | 34.0 | 37.5 | |
@@
SimLTD MixPL [6] | Objects365-unlabeled | DINO | R50 | 481 | 37.2 | 31.2 | 36.5 | 40.5 |
SimLTD MixPL [6] | COCO-unlabeled2017 | DINO | R50 | 477 | 37.7 | 32.4 | 36.9 | 40.9 |
@@ 3.4 Step 3: Fine-Tuning on
At this stage, we have two separate models with a shared representation, one optimized on head classes and the other on tail classes. We wish to unify the two models into one for efficient single-pass inference on test samples containing both head and tail categories, analogous to previous research on few-shot detection [38, 35]. Algorithm 1 provides the pseudocode for the merging scheme referenced in Figure 3.
Note that we merge parameters at the detector module and reuse the pre-trained head representations for the rest of the network. We fine-tune the unified detector on composed of instances, or shots, per class sampled from the long-tailed training set. We update only the box classifier and regressor with a reduced learning rate to slowly adapt them to tail classes while preserving the pre-trained accuracy on head classes. Similar to class-incremental learning, we include both head and tail classes in for exemplar replay to avoid catastrophic forgetting. We form via random sampling whereas prior work resorted to a complicated scheme of confidence-guided exemplar replay [12].
@@ 4.1 Evaluation Protocol
We benchmark our approach on the challenging LVIS v1 dataset, which has training and validation images over classes. We use the standard LVIS evaluator to compute the detection metric mAP for all classes, without test-time augmentation, along with AP, AP, and AP for the rare, common, and frequent categories. We sample three times with different random seeds and report on the averaged metrics to capture statistical variability. Following prior studies, we focus on the performance gains of mAP and AP in our comparative analysis.
@@ 4.2 Implementation Details
We implement our models using PyTorch [29] and MMDetection [5], and train on 8 A6000 GPUs. We pre-train most models in Step 1 for 540K iterations, perform transfer learning in Step 2 for 20K iterations, and finally fine-tune for 80K iterations. One experiment takes between 8 hours and 15 days to complete, depending on the scope. Full details are given in the submitted code.
High-Quality Supervised Baseline We construct a stronger supervised baseline than previously by combining our multi-stage training recipe with diverse data augmentations. We explore various ResNet [16] backbones for both Faster R-CNN and DETR-based models. We also include FPN [24] for additional feature extraction with Faster R-CNN. The supervised baseline is important to our framework because it serves as the basis for the teacher model to generate reliable pseudo targets for semi-supervised learning.
@@ 4.3 Main Results
Table 2 reports on the effectiveness of our approach against existing methods representing the state of the art on LVIS. We make a conscientious effort to have a fair comparison across multiple dimensions of backbone, detector, and external data source since not all methods follow a single established experimental protocol. Our SimLTD supervised baseline with Faster R-CNN outperforms all methods using related detectors. The gains are convincing, with margins up to AP and AP. We observe a similar trend when comparing our supervised baseline using Deformable DETR; SimLTD is on par with RichSem on AP while surpassing the competition on AP, including RichSem, with a significant improvement by up to AP points.
For the methods requiring external data, our semi-supervised models also deliver impressive performance. When equipped with MixTeacher and Faster R-CNN, SimLTD exceeds the competition by up to AP while being competitive on AP. Furthermore, SimLTD scales well by coupling with MixPL and the DINO detector to achieve new state-of-the-art results from harnessing only unlabeled images. Interestingly, the performance difference between COCO and Objects365 unlabeled images is small, signifying that SimLTD can exploit meaningful pseudo-label supervision from a large uncurated database with a different distribution than the training dataset. Figure 6 visualizes SimLTD detections on select LVIS validation images.
Ancillary to the main LVIS results, we also evaluate on Objects365 to showcase the generality of our framework. Following NorCal [28], we sample 30% of the training set and split it into classes appearing in images and classes appearing in images. The evaluation is partitioned into four groups of categories based on the count of training images by category. Figure 7 illustrates the large difference of instance counts between LVIS and Objects365, with Objects365 having more labels than LVIS for its most abundant object. SimLTD outperforms the existing baseline across the board in the validation results. We fine-tune SimLTD with shots, a parameter that changes by dataset. Section 4.5 gives a detailed ablation analysis regarding the impact on AP from varying the number of shots for fine-tuning with .
@@ 4.4 Comparison to the State of the Art
SimLTD vs. CascadeMatch [44] Although both SimLTD and CascadeMatch leverage COCO-unlabeled2017 for semi-supervised LTD, there are major differences between the two. First, CascadeMatch is trained end-to-end in a single stage whereas we take the decoupled approach. Second, CascadeMatch adopts the stronger Cascade R-CNN [2] compared to Faster R-CNN in SimLTD. CascadeMatch follows the AP protocol [8], which replaces the standard maximum 300 detections per image by a cap of 10K detections per class from the entire validation set. Despite the disadvantage of a simpler model, Table 3 shows that SimLTD outperforms CascadeMatch by notable margins in almost every measure. These superior results lend further support to the merit of our multi-stage training strategy.
SimLTD vs. Dong et al. [12] on Multi-Stage Learning We compare our multi-stage training strategy to that of Dong et al., which also utilizes a three-step procedure of pre-training followed by fine-tuning and knowledge distillation. We focus our analysis on their powerful Deformable DETR model, which yields similar results to our simpler Faster R-CNN model. When we train with the same capable Deformable DETR architecture, Table 2 shows that SimLTD exceeds their model by outsized margins of AP and AP. These remarkable gains are directly attributed to our multi-stage learning approach, which is carefully designed to optimize for accuracy on both head and tail classes.
SimLTD vs. RichSem [27] on Using Extra Data Recall that RichSem relies on the CLIP classifier (pre-trained on 400M image-caption pairs) and image-supervision from an additional 1.5M images to produce the state-of-the-art results reported in Table 2. However, such strict dependencies are infeasible when the method is applied to a bespoke dataset outside of generic objects. By contrast, our SimLTD leverages optional unlabeled images to deliver better performance than RichSem without resorting to either CLIP or auxiliary ImageNet labels. When we remove external data, SimLTD substantially outperforms RichSem by AP in the fully supervised setting, implying that the success of RichSem is sensitive to the contributions of CLIP and ImageNet supervision. Our framework carries the benefit of being robust across settings with and without external data.
@@
Multi-Stage Baseline (w/ Random Resize) | 26.8 | 17.9 |
Multi-Stage Baseline +Photometric Jittering | 26.9 | 18.2 |
Multi-Stage Baseline ++Repeat Sampling | 27.1 | 19.3 |
Multi-Stage Baseline +++Copy-Paste (Ours) | 29.0 | 20.9 |

### v2

@@ Abstract
While modern visual recognition systems have made significant advancements, many continue to struggle with the open problem of learning from few exemplars. This paper focuses on the task of object detection in the setting where object classes follow a natural long-tailed distribution. Existing methods for long-tailed detection resort to external ImageNet labels to augment the low-shot training instances. However, such dependency on a large labeled database has limited utility in practical scenarios. We propose a versatile and scalable approach to leverage optional unlabeled images, which are easy to collect without the burden of human annotations. Our SimLTD framework is straightforward and intuitive, and consists of three simple steps: (1) pre-training on abundant head classes; (2) transfer learning on scarce tail classes; and (3) fine-tuning on a sampled set of both head and tail classes. Our approach can be viewed as an improved head-to-tail model transfer paradigm without the added complexities of meta-learning or knowledge distillation, as was required in past research. By harnessing supplementary unlabeled images, without extra image labels, SimLTD establishes new record results on the challenging LVIS v1 benchmark across both supervised and semi-supervised settings.
@@ 1 Introduction
Another interesting direction is to leverage external data to augment the training instances in LVIS. The intuition is that while the rare objects may not appear frequently in natural scenes, they can be found in abundance from object-centric sources such as ImageNet [9]. By training on the union of scene-centric LVIS and object-centric ImageNet images, these weakly supervised methods [45, 49, 27] overcome the skewed class distribution by enriching the tail instances with additional whole-image labels. Although appealing, this approach places a hard requirement on having access to a large database of 14M labeled images across 21K object classes, an assumption not necessarily suitable for many practical settings outside of LVIS.
@@
We propose a simple and scalable framework, aptly named SimLTD, to answer this question. We deconstruct the LTD problem into three stages consisting of (i) pre-training on head classes, (ii) transfer learning on tail classes, and (iii) fine-tuning on a small sampled set composed of both head and tail classes. Unlike previous research on multi-stage training [19, 12], our framework allows for the optional use of unlabeled images to further boost LTD. We learn with supplementary unlabeled data in a semi-supervised manner via pseudo-labeling, and therefore do not explicitly rely on any additional instance-level or image-level supervision.
Our work addresses several challenges associated with the head-to-tail model transfer paradigm [18, 39]. For one, we find that the vanilla transfer of data-rich head representations to data-scarce tail classes does not yield sufficient performance improvement because the distribution of head classes is also skewed. Moreover, the naïve application of unlabeled data can aggravate the model’s inductive bias because the unlabeled samples may follow a similar long-tailed distribution. In such case, the trained model will tend to generate more pseudo labels for the head classes, resulting in a larger degree of pseudo-class imbalance. We overcome these obstacles by incorporating data augmentations specifically designed to mitigate class imbalances in both head and tail training stages. Extensive experiments in §4.3 show that stronger pre-trained models on head classes, with and without unlabeled images, transfer well to the tail classes, leading to a desirable solution for long-tailed detection.
@@ 2 Related Work
Leveraging Extra Data for LTD There is an emerging trend of using the abundant weak-supervision of ImageNet labels to solve the LVIS problem. Although LVIS is a detection dataset of natural scenes and ImageNet is a classification dataset of object images, they both share 997 overlapping classes between their vocabularies, the intersection of which provides a rich source of 1.5M extra images to sample additional LVIS instances. To effectively learn from a mixed LVIS-ImageNet dataset with a domain gap, MosaicOS [45] leverages mosaic augmentation [1], whereas Detic [49] and RichSem [27] rely on the CLIP [30] classifier to map the semantic concepts between LVIS and ImageNet targets.
@@ 3 The SimLTD Framework
As illustrated in Figure 2, given a long-tailed dataset with categories, we split it into two disjoint subsets: with frequent and common categories appearing in images and with rare classes appearing in images. Furthermore, we have access to an unlabeled source of unknown class distribution. Our goal is to use a combination of labeled and unlabeled images to train a unified model optimized for accuracy on a test set comprising both classes in .
Our SimLTD framework consists of three easy steps: (i) representation learning on , (ii) transfer learning on , and (iii) fine-tuning on , a reduced dataset composed of instances per class randomly sampled from . Note that , , and are all still imbalanced, but not as severe as the original long-tailed . We leverage optional unlabeled images in both Steps 1 and 2 but do not explicitly need them for effective LTD. Indeed, experiments in §4.3 show that our fully supervised baselines exhibit excellent performance and scalability even without unlabeled data. A diagram of SimLTD is depicted in Figure 3.
@@ 3.1 The Devil is in the
Discussion We observe a clear trend indicating stronger pre-trained representations, as measured by the AP score on COCO, generally lead to improved rare class detection. This result is both intriguing and encouraging since the models were pre-trained on COCO, a dataset of different scope and size than LVIS. Training from scratch brings out the worst performance with half of the accuracy. The implication of this simple experiment is two-fold: (1) we corroborate prior studies by showing low-shot learning can be improved with transferred representations; and (2) our framework opens opportunities to self-supervised, semi-supervised, and multi-modal learning, all of which have demonstrated significantly better performance than supervised pre-training. Motivated by these insights, we propose to learn powerful representations on and transfer them to for long-tailed detection. While the past attempts at head-to-tail model transfer [39, 19, 12] could only work by incorporating an extra module for meta-learning or knowledge distillation, we now describe our three-step approach to accomplish this goal without the unnecessary complexities.
@@ 3.2 Step 1: Representation Learning on
As summarized in Table 1, we leverage well-known data augmentations to train , which include random image resizing, horizontal flipping, and photometric distortion. We also combine Simple Copy-Paste (SCP) [14] together with Repeat Factor Sampling (RFS) [15] to combat the class imbalance in at both the image and instance levels, analogous to prior research on the importance of image and instance resampling for LTD [37, 42]. Taking a step further to learn even stronger representations on , we propose to train in a semi-supervised manner with unlabeled images by way of pseudo-labeling. We explore three successful methods of increasing effectiveness to advance semi-supervised representation learning on : SoftER Teacher [35], MixTeacher [25], and MixPL [7].
@@
| Seesaw Loss [CVPR’21] [36] | N/A | Faster R-CNN | R101-FPN | – | 27.8 | 18.7 | 27.0 | 32.8 |
@@
| SimLTD Supervised | Faster R-CNN | R101-FPN | 83 | 31.7 | 24.3 | 32.1 | 34.6 | |
| SimLTD Supervised | Deformable DETR | R101 | 234 | 37.0 | 31.9 | 36.6 | 39.6 | |
@@
| SimLTD Supervised | Faster R-CNN | R50-FPN | 81 | 29.0 | 20.9 | 29.7 | 31.9 | |
| SimLTD Supervised | Deformable DETR | R50 | 215 | 35.0 | 32.0 | 34.0 | 37.5 | |
| RichSem [NeurIPS’23] [27] | N/A | DINO | Swin-T | – | 38.8 | 30.8 | 36.4 | 45.0 |
| SimLTD Supervised | DINO | Swin-T | 310 | 41.1 | 33.6 | 40.1 | 45.4 | |
| DiffusionDet [ICCV’23] [6] | 4 @ 300 | Swin-B | – | 42.0 | 34.8 | 40.9 | 46.4 | |
| RichSem [NeurIPS’23] [27] | DINO | Swin-B | – | 46.4 | 38.5 | 45.1 | 51.3 | |
| SimLTD Supervised | DINO | Swin-B | 414 | 47.2 | 42.7 | 46.7 | 49.9 | |
| RichSem [NeurIPS’23] [27] | DINO | Swin-L | – | 49.7 | 42.8 | 49.2 | 53.4 | |
| SimLTD Supervised | DINO | Swin-L | 460 | 49.8 | 42.4 | 50.4 | 52.4 | |
@@
| SimLTD MixPL [7] | Objects365-unlabeled | DINO | R50 | 447 | 38.3 | 30.8 | 37.5 | 42.6 |
| SimLTD MixPL [7] | COCO-unlabeled2017 | DINO | R50 | 446 | 38.5 | 31.6 | 37.5 | 42.5 |
| RichSem [NeurIPS’23] [27] | ImageNet-21K Labels | DINO CLIP | Swin-T | – | 41.6 | 37.3 | 39.7 | 45.5 |
| SimLTD MixPL [7] | COCO-unlabeled2017 | DINO | Swin-T | 482 | 42.5 | 35.7 | 42.4 | 45.6 |
| RichSem [NeurIPS’23] [27] | ImageNet-21K Labels | DINO CLIP | Swin-B | – | 48.2 | 46.5 | 46.5 | 51.0 |
| SimLTD MixPL [7] | COCO-unlabeled2017 | DINO | Swin-B | 794 | 49.0 | 43.4 | 49.0 | 51.5 |
| RichSem [NeurIPS’23] [27] | ImageNet-21K Labels | DINO CLIP | Swin-L | – | 52.0 | 50.2 | 51.5 | 53.3 |
| SimLTD MixPL [7] | COCO-unlabeled2017 | DINO | Swin-L | 1168 | 51.5 | 45.0 | 52.4 | 53.3 |
@@ 3.4 Step 3: Fine-Tuning on
At this stage, we have two separate models with a shared representation, one optimized on head classes and the other on tail classes. We wish to unify the two models into one for efficient single-pass inference on test samples containing both head and tail categories. Algorithm 1 provides the pseudocode for the merging scheme referenced in Figure 3.
Note that we merge parameters at the detector module and reuse the pre-trained head representations for the rest of the network. We fine-tune the unified detector on composed of instances, or shots, per class sampled from the long-tailed training set. We update only the box classifier and regressor with a reduced learning rate to slowly adapt them to tail classes while preserving the pre-trained accuracy on head classes. We form to include both head and tail classes for exemplar replay to avoid catastrophic forgetting.
@@ 4.1 Evaluation Protocol
We benchmark our approach on the challenging LVIS v1 dataset, which has training and validation images over classes. We compute the detection metric mAP for all categories, without test-time augmentation, along with AP, AP, and AP for the rare, common, and frequent classes. We sample three times with different random seeds and report on the averaged metrics to capture statistical variability. Following prior studies, we focus on the gains of mAP and AP in our comparative analysis.
@@ 4.2 Implementation Details
We implement our models using PyTorch [29] and MMDetection [5], and train on 8 A6000 GPUs. We pre-train in Step 1 for up to 540K iterations, perform transfer learning in Step 2 for 20K iterations, and fine-tune for 80K iterations. See our open-source code for the full reproducible details.
High-Quality Supervised Baseline We construct a strong supervised baseline by combining our training recipe with diverse data augmentations. We explore various ResNet [16] and Swin [26] backbones, along with FPN [23], for feature extraction. The supervised baseline is important to our framework because it serves as the basis for the teacher model to propose reliable pseudo targets for semi-supervised learning.
@@ 4.3 Main Results
Table 2 reports on the effectiveness of our approach against existing methods representing the state of the art on LVIS. Our SimLTD supervised baseline with Faster R-CNN outperforms all methods using related detectors. The gains are convincing, with margins up to AP and AP. We observe a similar trend when comparing our supervised baseline using DETR-based models. SimLTD demonstrates compelling performance and scalability across a multitude of backbones and detectors without the need for extra data.
For the methods requiring external data, our semi-supervised models also deliver impressive performance. When equipped with MixTeacher and Faster R-CNN, SimLTD exceeds the competition by up to AP while being competitive on AP. Furthermore, SimLTD scales well by coupling with MixPL and transformer-based models to achieve new state-of-the-art results from harnessing only unlabeled images. SimLTD works equally well with both COCO and Objects365 unlabeled images, signifying that our model can extract meaningful pseudo-label supervision from a large uncurated database with a distribution different from the training dataset. Figure 6 visualizes example SimLTD detections on select LVIS validation images.
Ancillary to the main LVIS results, we also evaluate on Objects365 to showcase the generality of our framework. Following NorCal [28], we sample 30% of the training set and split it into classes appearing in images and classes appearing in images. SimLTD outperforms the existing baseline across the board in Figure 7. We fine-tune SimLTD with shots, a parameter that changes by dataset. Section 4.5 gives a detailed ablation analysis regarding the impact on AP from varying the number of shots for fine-tuning with .
@@ 4.4 Comparison to the State of the Art
SimLTD vs. CascadeMatch [44]
Although both SimLTD and CascadeMatch leverage COCO-unlabeled2017 for semi-supervised LTD, there are major differences between the two. First, CascadeMatch is trained end-to-end in a single stage, whereas we take the decoupled approach. Second, CascadeMatch adopts the stronger Cascade R-CNN [2] compared to Faster R-CNN in SimLTD. CascadeMatch follows the AP protocol [8], which replaces the standard maximum 300 detections per image by a cap of 10K detections per class from the entire validation set. Despite the disadvantage of a simpler model, Table 3 shows that SimLTD outperforms CascadeMatch by notable margins in almost every measure. These superior results lend further support to the merit of our multi-stage training strategy.
SimLTD vs. Dong et al. [12] on Multi-Stage Learning We compare our multi-stage training strategy to that of Dong et al., which also utilizes a three-step procedure of pre-training followed by fine-tuning and knowledge distillation. We focus our analysis on their powerful Deformable DETR model, which yields analogous results to our simpler Faster R-CNN model. When we train with the same capable Deformable DETR architecture, Table 2 shows that SimLTD exceeds their model by outsized margins of AP and AP. These remarkable gains are directly attributed to our multi-stage learning approach, which is carefully designed to optimize for accuracy on both head and tail classes.
SimLTD vs. RichSem [27] on Using Extra Data Recall that RichSem relies on the CLIP classifier (pre-trained on 400M image-caption pairs) and image-supervision from an additional 1.5M images to produce the state-of-the-art results reported in Table 2. However, such strict dependencies are impractical when the method is applied to a bespoke dataset outside of generic objects. By contrast, our SimLTD leverages unlabeled images, without resorting to either CLIP or auxiliary ImageNet labels, to deliver better results than RichSem with the ResNet backbone and comparable performance with the Swin backbones. When we remove external data, SimLTD substantially outperforms RichSem by AP in the fully supervised setting, implying that the success of RichSem is sensitive to the contributions of CLIP and ImageNet supervision. Our framework carries the benefit of being robust across settings with and without external data.
@@
| Multi-Stage Training (w/ Random Resize) | 26.8 | 17.9 |
| Multi-Stage Training +Photometric Jittering | 26.9 | 18.2 |
| Multi-Stage Training ++Repeat Sampling | 27.1 | 19.3 |
| Multi-Stage Training +++Copy-Paste (Ours) | 29.0 | 20.9 |